Calculate the vapour pressure of a solvent above a solution containing 83.99g of solute x(molecular weight 195.62) per 1200g of h20 at 20°c.the vapour pressure of h2O at 20°c is 17.53mmhg
1
Expert's answer
2020-11-09T13:50:33-0500

pH2O=pH2O**xH2O (Raoult's law)

nH2O=1200/18 = 66.667 mol

nx = 83.99/195.62 = 0.429

xH2O = 66.667/(66.667+0.429) = 0.99361

pH2O=17.53*0.99361 = 17.42 mmhg